The following information applies to the next two questions Dée Trader opens a brokerage account and purchases 300 shares of Internet Dreams at $40 per share. She borrows $4,000 from her broker to help pay for the purchase. Question 1 (1 point) What is the initial margin in Dées account when she first purchases the stock? Enter your answer as a decimal, rounded to two decimal places. Your Answer Answer Question 2 (1 point) If the share price falls to $30 per share, what is the margin in her account? Enter your answer as a decimal, rounded to two decimal places. Your Answer: Answer
Answer #1

(1) Number of Shares Bought = 300 and Price per Share = $ 40

Total Value of Shares Purchased = 300 x 40 = $ 12000

Amount Borrowed = $ 4000

Initial Margin = Buyer's Own Contribution = Buyer's Equity = 12000 - 4000 = $ 8000

(2) Current Share Price = $ 30

Total Value of Shares = 300 x 30 = $ 9000

Amount Borrowed = $ 4000

Remaining Margin = 9000 - 4000 = $ 5000